solve by the elimination method
3x-2y=7
-x+5y=2

Answer by oberobic(2304)

3x - 2y = 7
-x + 5y = 2
.
multiply the 2nd equation by 3 and add
.
3x - 2y = 7
-3x +15y = 6
--------------
0x + 13y = 13
.
13y = 13
.
y = 1
.
substitute y=1 to find 'x'
.
3x -2y = 7
3x - 2(1) = 7
3x - 2 = 7
3x = 9
x = 3
.
always check your work, so substitute these values into the 2nd equation to compare to the known answer
.
-(3) + 5(1) = 2
correct.
.
Answer: x=3, y = 1
.
Done.